OpenStudy (toribyrd33):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 3?

OpenStudy (toribyrd33):

an = 5 − 3(n − 1); all integers where n ≥ 1 an = 5 − 3(n − 1); all integers where n ≥ 0 an = 5 − 2(n − 1); all integers where n ≥ 0 an = 5 − 2(n − 1); all integers where n ≥ 1

OpenStudy (welshfella):

general form = a + d(n - 1) where a = first term and d = common difference d = second term - first

OpenStudy (welshfella):

n>= 1

OpenStudy (toribyrd33):

is it C?

OpenStudy (toribyrd33):

@welshfella

OpenStudy (welshfella):

sorry - i keep on get disconnected

OpenStudy (welshfella):

no because n >= 1

OpenStudy (toribyrd33):

did you mean\[n \ge1\]

OpenStudy (welshfella):

yes

OpenStudy (toribyrd33):

oh ok

OpenStudy (toribyrd33):

thank you

OpenStudy (welshfella):

yw